# Gianluca Monaci          Rafael S. Rezende          Romain Deffayet          Gabriela Csurka Guillaume Bono          Hervé Déjean          Stéphane Clinchant          Christian Wolf

### v1

@@ 3 Retrieval-augmented agent
In summary, the agent receives at each timestep a new visual observation $\mathbf{x}_{t}$, which it encodes, compares with the goal $\mathbf{g}$, and enriches with additional context information retrieved from the database $\mathcal{D}$. This representation is then passed to a recurrent policy. See Fig. 3 for a visual overview.
@@ 3.1 Trainable context encoder $c$
$\mathbf{e}_{t,n}=\leavevmode\nobreak\ \leavevmode\nobreak\ d(\mathbf{x}_{t},\mathbf{r}_{t,n})$ // ObjNav (5)
$\mathbf{e}_{t,n}=[d(\mathbf{x}_{t},\mathbf{r}_{t,n})\ d(\mathbf{g},\mathbf{r}_{t,n})].$ // ObjNav or ImageNav (6)
At this point, we obtain a list of embeddings $\mathbf{e}_{t,n}$ that benefit from the rich inductive bias provided by the geometric FM, including information about pose and visibility of the context image $\mathbf{r}_{t,n}$ from the observed image $\mathbf{x}_{t}$ and goal image $\mathbf{g}$. We must now derive a single context embedding $\tilde{\mathbf{c}}_{t}=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})$ that provides useful information to the agent. We opt for two variants:
Variant 1: Gumbel soft-max selector – The agent selects a single embedding from the list by predicting a distribution over items in the retrieved list followed by a Gumbel soft-max sampler [37]. More precisely,
@@
where $\sigma$ is the soft-max distribution over items in the list. We perform the sampling and the computation of gradients using the Gumbel soft-max trick [37]. We motivate this choice from the pre-training of the geometric FM $d$ in Eq. 5 and Eq. 6: the latent vector $z=d(a,b)$ has been pre-trained to provide overlap information between images $a$ and $b$, which we conjecture to be highly correlated with the relevance of the corresponding context item. Additionally, since the weights $\alpha_{i}$ are calculated independently, the model can be zero-shot applied to any context size at test time.
Variant 2: attention-based encoder – A transformer-based model is potentially capable of integrating multiple context tokens $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ into a single fused embedding, effectively changing and enriching the original embedding space. We use a transformer encoder network with 2 layers of 4-heads self-attention and predict the output by concatenating the context tokens and processing them with a 2-layer MLP.
@@ 4 Retrieval mechanisms for RANa
In all tasks and use cases, the retrieval process consists of a nearest neighbor search on representations obtained with pre-trained models. Depending on the navigation task at hand, we either use DINOv2 [55] when we have a visual query or Open-CLIP [35] for textual queries. For zero-shot use-cases that require replacing the task goal $\mathbf{g}$ with a retrieved image goal $\mathbf{g}^{\mathcal{R}}$ ([Uncaptioned image], [Uncaptioned image]), we set $\mathbf{g}^{\mathcal{R}}$ as the nearest neighbor of $\mathbf{g}$ from the dataset $\mathcal{D}$. For use-cases that require building a context set to be passed to the encoder $c$ ([Uncaptioned image], [Uncaptioned image], [Uncaptioned image]), we investigate two strategies: (i) static context, where only the goal is used to build a fixed context kept throughout the episode, and (ii) dynamic context, which also depends on the current observation and is recomputed at each timestep.
@@ 4.1 Static context
We build the static context by ranking images from the dataset based on DINOv2 or CLIP feature similarity with the goal and selecting the top-$N$ nearest neighbors.
Diversity – providing multiple near-duplicates to the selector might not be very helpful. Moreover, in ObjectNav a scene may contain several instances of the target class and it is better to provide context images showing different instances. We therefore resort to Maximum Margin Relevance (MMR) [11] re-ranking to increase diversity. Assuming a relevance score vector $\omega$ (DINOv2 or CLIP similarities) providing a ranked shortlist (top-100 in our experiments), and a matrix $\Omega$ encoding the similarity for each pair of images in the ranked shortlist, MMR greedily selects at each step (rank) $i$ the element $r_{i}$ that maximizes the re-ranking criterion:
@@
where $\beta\in[0,1]$ is a mixture parameter (set to $0.5$ in all our experiments) and $\mathcal{S}_{i-1}$ is the set of images already selected. We then select the top-$N$ highest MMR.
CLIP similarity normalization – While CLIP allows to easily compare any open-vocabulary object category with retrieval gallery images, it has been observed in the context of zero-shot image classification that directly ranking images by similarity to the category is suboptimal [58]. To address this issue for ObjectNav, we leverage the fact that we often target a pre-defined set of classes $\mathcal{O}$ (e.g. 6 in HM3D), and rescale the score of each image in $\mathcal{D}$ via softmax:
@@ 4.2 Dynamic context set
It can be interesting to build the context using not only the goal but also the current observation, for example to find intermediate waypoints between the current observation and the goal. We propose to use an affinity graph based on similarities of images in the dataset $\mathcal{D}$ to find such waypoints.
DINOv2-Graph – Given $\mathcal{D}$, we can build the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between all pairs of images in $\mathcal{D}$ (we use DINOv2 [55] features). Using $\Omega_{\mathcal{D}}$, we can build a graph whose nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity. Then, we find in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively. Finally, we populate the context $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})$ with $r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$ and images sampled on the shortest path in the graph between them. We provide more details about the graph and its relation to topological maps in the Supplementary.
@@ 5 Experimental results
We train and evaluate our agents on the Habitat simulator and platform [64] according to the standard ImageNav, Instance-ImageNav and ObjectNav task definitions.
@@
Retrieval-augmented navigation benchmarks – There is no publicly available navigation benchmark that provides a database of observations that can be exploited during training and evaluation. We thus propose to derive ImageNav and ObjectNav benchmarks from standard Habitat tasks [64], modifying them as little as possible to enable training and evaluating retrieval-augmented agents. For ImageNav, we use the Gibson dataset [77], consisting of 72 train and 14 eval scenes, and for ObjectNav, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 [60, 81]. All agents use only one $112{\times}112$ RGB sensor. For each scene, unless otherwise stated, we generate a retrieval set $\mathcal{D}$ of $1,000$ FPV images by letting agents navigate the environment and store images in $\mathcal{D}$. During training and evaluation, the retrieval-augmented agent can select dataset images from the scene it operates in. Further details on dataset creation are in Supplementary.
@@
Implementation details – For ImageNav ([Uncaptioned image], [Uncaptioned image]), we use the experimental setup for Gibson dataset [68] where the goal is a random view taken by the agent camera, as in, e.g.[48, 52, 6]. We base our agents on the DEBiT architecture [6] by starting from the official codebase and weights provided by the authors [^2] https://github.com/naver/debit, and extend them as described in Sec. 3. We consider two DEBiT backbones, base (DEBiT-b) with a binocular encoder with 55M parameters, and tiny (DEBiT-t), with a 17M-parameters encoder to allow faster execution for the ablation studies on Sec. 5.4. The corresponding retrieval-augmented models are RANa-b and RANa-t. In Sec. 5.2 we use the same agent and weights in zero-shot fashion for Instance-ImageNav and ObjectNav when the task goal is replaced with a retrieved image goal ([Uncaptioned image], [Uncaptioned image]).
For the ObjectNav base architecture with a textual goal ([Uncaptioned image], [Uncaptioned image]), we use the default experimental setup of HM3DSem-v0.2 [60, 10]. We modify the RGB observations to be $112{\times}112$ instead of $480{\times}640$ as in the original ObjectNav definition, so we can load the pretrained weights of the hwRN18 visual encoder $x$ from DEBiT-t [6]. We use the OpenCLIP [35] model trained on Laion-2B [65] [^3] https://github.com/mlfoundations/open_clip/.
Model Retrieval Encoder SR SPL
[Uncaptioned image] DEBiT-t [6] 79.9 51.4
[Uncaptioned image] RANa-t DINOv2 Gumbel 83.0 52.3
@@
[Uncaptioned image] RANa-t DINOv2-Graph Attention 84.7 63.7
[Uncaptioned image] DEBiT-b [6] 83.4 56.8
[Uncaptioned image] RANa-b DINOv2 Gumbel 88.3 57.6
@@
[Uncaptioned image] RANa-b DINOv2-Graph Attention 90.7 71.8
[Uncaptioned image] hwRN18+CLIP 42.9 23.0
[Uncaptioned image] RANa-t OpenCLIP Gumbel 52.1 26.9
[Uncaptioned image] RANa-t OpenCLIP Attention 49.7 26.0
Table 2: Effect of retrieving context on ImageNav and ObjectNav. models finetuned for the same number of steps as corresponding RANa models for fairness of comparison.
Model training – All models, retrieval-augmented or not, are trained with PPO up to 200M steps. Unless specified otherwise, the geometric FM $d$ and encoders $x$ and $l$ are loaded from DEBiT and kept frozen, as are OpenCLIP and DINOv2. For retrieval-augmented agents, the context encoder $c$ and policy network $\pi$ are learned from scratch. Compared to DEBiT and hwRN18+CLIP , the GRU memory of RANa receives the additional context input $\tilde{\mathbf{c}}_{t}$, therefore we expand the DEBiT/hwRN18+CLIP GRU input matrix to reach the required dimension and finetune it. Further details and ablations of this choice are reported in Supplementary. The reward definition is inspired by PointGoal [15] and ImageNav [6] rewards and given as $r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$ where $K{=}10$, $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal, and slack cost $\lambda{=}0.01$ encourages efficiency.
Retrieval – We consider two context encoding approaches, Gumbel and Attention, and three retrieval mechanisms based on DINOv2 and DINOv2-graph for ImageNav, and OpenCLIP for ObjectNav tasks. Unless otherwise specified, we use a context size $N{=}8$. Results for ObjectNav use a retrieval dataset of size $10$k images (see Sec. 5.4 and Supplementary for ablations).
Metrics – Navigation performance is evaluated by the success rate (SR), i.e., the percentage of episodes terminated within a distance of ${<}1$m from the goal after the agent has called the STOP action, and SPL [2], i.e., SR weighted by the optimality of the path: $\textit{SPL}=\frac{100}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$ where $S_{i}$ is a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
5.1 Retrieval improves navigation
@@
Tab. 2 shows results of the proposed RANa agents with retrieved context on ImageNav and ObjectNav tasks, which exhibit significant performance improvements across all settings compared to the DEBiT baseline and the ObjectNav baseline models [^4] To make baselines comparable, the loaded checkpoints continued training for the same additional 200M steps updating GRU and policy only.. For static context retrieval (both using DINOv2 for ImageNav or OpenCLIP for ObjectNav) the Gumbel encoder shows better results, arguably because there is no inherent ordering in the context and the retrieved CroCo-based features contain strong information about the relative pose between context elements and observation. The attention-based encoder works better for the dynamic graph-based context, where reasoning about the order of context elements becomes important. The graph-based context combined with attention obtains the best results, with a huge boost in SPL ($+14.2$ and $+12.3$ points) compared to the DEBiT-b and DEBiT-t baselines, respectively. For all ObjectNav runs, we found that finetuning $x$ while training on HM3D improved performance, as the visual encoder of DEBiT was only trained on Gibson observations. Moreover, since ImageNav and ObjectNav have different action spaces, we learn the action embedding $l$ from scratch.
Model SR SPL Binocular encoder
ZSON [47] 36.9 28.0 Observation finetuned, goal frozen
OVRL [80] 54.2 27.0 Finetuned
VC1-ViT-L [49] 81.6 - Finetuned
OVRL-v2 [79] 82.0 58.7 Finetuned
DEBiT-b no adapters [6] 83.0 55.6 Frozen
DEBiT-L no adapters [6] 82.0 59.6 Frozen
DEBiT-L+adapters[6] 94.0 71.7 Finetuned with adapters
RANa-b Graph-Attn 90.7 71.8 Frozen (base DEBiT-b no adapters)
Table 3: ImageNav state-of-the-art results. these models use additional GPS+Compass input providing exact localization.
Comparison to ImageNav SotA – In Tab. 3 we compare RANa ImageNav with state-of-the-art (SotA) models that typically finetune (with a navigation loss) visual encoders based on CLIP [47], ViT [80, 49, 79] and geometric FM [6] used for observation-goal comparison. RANa boosts the performance of DEBiT-b, which uses frozen binocular encoders, and achieves performance on par with SotA method DEBiT-Large [6] finetuned with adapters [17], which use an involving adaptation process to tune visual encoders to the task. In contrast, we achieve strong performance through (i) scalable context encoding and (ii) a generic frozen visual FM, enabling several zero-shot tasks (Sec. 5.2).
Comparison to ObjectNav SotATab. 4 compares the results of RANa ObjectNav tiny model with competing approaches. Compared to methods of similar sensor settings (one FPV RGB camera), RANa-t outperforms previous SotA by $+9.7$ on SR and $+7.7$ on SPL. Neither ZSON [47] nor PSL [70] were trained on ObjectNav episodes, but on HM3D ImageNav episodes, which might give a small advantage to RANa. Hence we also compare RANa-t to methods that use extra sensors, such as depth camera, GPS+compass, or 6 camera sensors, as well as LLM-guided navigation. Here RANa-t shows competitive results, outperforming several strong methods such as ESC [86] and PixNav [10]. Compared to the considerably more complex L3MVN [83], which leverages depth and pose to construct a semantic map and uses GPT-2 to guide exploration via object searching, RANa-t is competitive, being $-2.1$ below on SR but $+1.4$ above on SPL, meaning that the retrieved context guides our agent to more efficient routes to object goals.
@@
100 dataset images suffice to improve performanceTab. 5 shows that the size of the retrieval database can be small, $100$ is enough for the considered Gibson scenes in ImageNav. Additional experiments for ObjectNav and zero-shot tasks are presented in Supplementary.
Model Extra sensors w/ map w/ LLM SR SPL
PixNav [10] Pano RGB 37.9 20.5
ESC [86] Depth, odom 39.2 22.3
L3MVN [83] Depth, odom 54.2 25.5
CoW [27] Depth, odom 6.1 3.9
GoW [27] Depth, odom 32.0 18.1
ProcTHOR [19] Depth, odom 13.2 7.7
ZSON [47] 25.5 12.6
PSL [70] 42.4 19.2
RANa-t (Ours) 52.1 26.9
Table 4: Comparison to ObjectNav state-of-the-art methods.
100 1k 10k 20k
@@
Table 5: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for ImageNav with retrieval using different database sizes, combinations of train and test conditions.

### v2

@@ 3 Retrieval-augmented agent
In summary, the agent receives at each timestep a new visual observation $\mathbf{x}_{t}$, which it encodes, compares with the goal $\mathbf{g}$, and enriches with additional context information retrieved from the database $\mathcal{D}$. This representation is then passed to a recurrent policy. See Fig.˜2 for a visual overview.
@@ 3.1 Trainable context encoder $c$
$\mathbf{e}_{t,n}=[g(\mathbf{x}_{t},\mathbf{r}_{t,n})\ g(\mathbf{g},\mathbf{r}_{t,n})].$ (3)
At this point, we obtain a list of embeddings $\mathbf{e}_{t,n}$ that benefit from the rich inductive bias provided by the geometric FM, including information about pose and visibility of the context image $\mathbf{r}_{t,n}$ from the observed image $\mathbf{x}_{t}$ and goal image $\mathbf{g}$. The list is compressed into a single context embedding $\tilde{\mathbf{c}}_{t}=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})$ that provides useful information to the agent, and for this we explore two variants.
Variant 1: Gumbel soft-max selector – The agent selects a single embedding from the list by predicting a distribution over items in the retrieved list followed by a Gumbel soft-max sampler (Jang et al., 2017):
@@
where $\sigma$ is the soft-max distribution over items in the list. We perform the sampling and the computation of gradients using the Gumbel soft-max trick (Jang et al., 2017). We motivate this choice from the pre-training of the geometric FM $g$ in Eq.˜3: the latent vector $z=g(a,b)$ has been pre-trained to provide overlap information between images $a$ and $b$, which we conjecture to be highly correlated with the relevance of the corresponding context item. Additionally, since the weights $\alpha_{i}$ are calculated independently, the model can be zero-shot applied to any context size at test time.
Variant 2: Attention-based encoder – A transformer-based model is potentially capable of integrating multiple context tokens $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ into a single fused embedding, effectively changing and enriching the original embedding space. We use a transformer encoder layer with 2 layers of 4-heads self-attention, which receives as input the $N$ context tokens summed with positional embeddings, and predicts the context feature $\tilde{\mathbf{c}}_{t}$ by concatenating the output tokens and feeding them to a 2-layer MLP.
@@ 4 Retrieval mechanisms for RANa
The retrieval process performs a nearest neighbor search on image representations obtained with DINOv2 (Oquab et al., 2024) resulting in a retrieved context. We investigate two context building strategies:
  1. 1.
    Static context, where only the goal is used to build a fixed context kept throughout the episode,
  2. 2.
    Dynamic context, which also depends on the current observation and is recomputed at each timestep.
@@ 4.1 Static context
[…]
Diversity – providing multiple near-duplicates to the selector might not be very helpful. We therefore resort to Maximum Margin Relevance (MMR) from Carbonell & Goldstein (1998) re-ranking to increase diversity. Assuming a relevance score vector $\omega$ (DINOv2 similarities) providing a ranked shortlist (top-100 in our experiments), and a matrix $\Omega$ encoding the similarity for each pair of images in the ranked shortlist, MMR greedily selects at each step (rank) $i$ the element $r_{i}$ that maximizes the re-ranking criterion:
@@
[…]
@@ 4.2 Dynamic context set
Dynamic context exploits the current observation at each time step, and we leverage this to use the context set to provide intermediate waypoint images to the augmented agent. Similar to topological maps we do this by building a graph and calculating shortest paths. However, in stark contrast to classical topological maps, (i) the graph is built from the images of the retrieval dataset only and can be pre-computed, (ii) no pose information is required, we use visual similarities only, and (iii) waypoint images are recommendations only, exploited or not by the agent through its retrieval context.
DINOv2-Graph – Given $\mathcal{D}$, we build the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between all pairs of images in $\mathcal{D}$ using DINOv2 feature cosine similarity (Oquab et al., 2024). From $\Omega_{\mathcal{D}}$ we can derive a graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, whose nodes $\mathcal{N}$ are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges $\mathcal{E}$ are weighted by similarity. Then, we find in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively. Finally, we populate the context $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})$ with $r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$ and images sampled on the shortest path in the graph between them.
Fig.˜3 shows three examples of shortest paths on the DINOv2-Graph, in green. For comparison only, we also visualize the shortest paths obtained by the ground-truth (GT) “Pose Graph”, not available to the agent. The path on the DINOv2-Graph well approximates the GT path. We provide more details about the graph construction and its relation to topological maps in Appendix˜C.
Figure 3: Shortest paths on DINOv2-Graph and Pose Graph - The most similar image $r_{1}^{\mathbf{x}_{t}}\in\mathcal{D}$ to the goal is highlighted in red and localized with a red star on the floorplan, while $r_{1}^{\mathbf{g}}\in\mathcal{D}$, in blue, is the most similar image to the current observation. Green crosses connects images along the shortest-path in the DINOv2-Graph (corresponding images in green) and purple crosses refer to poses on the Pose Graph, built using distances between GT camera poses, with corresponding images highlighted in purple at the bottom of the map.
Retrieving the goal – independent of the creation of a retrieval context, static or dynamic, in zero-shot settings we also use retrieval to replace goal $\mathbf{g}$ with a retrieved image goal $\mathbf{g}^{\mathcal{R}}$ ([Uncaptioned image], [Uncaptioned image]), setting $\mathbf{g}^{\mathcal{R}}$ as the nearest neighbor of $\mathbf{g}$ from the dataset $\mathcal{D}$.
@@ 5 Experimental results
We train and evaluate our agents on the Habitat simulator and platform (Savva et al., 2019) according to the standard ImageNav, Instance-ImageNav and ObjectNav task definitions.
@@
Retrieval-augmented navigation benchmarks – There is no publicly available benchmark that provides a database of observations that can be exploited during training and evaluation. We propose to derive our benchmarks from standard Habitat tasks (Savva et al., 2019), modifying them as little as possible to enable training and evaluating retrieval-augmented agents. We use the Gibson dataset (Xia et al., 2018), consisting of 72 train and 14 eval scenes, and for each scene, unless otherwise stated, we generate a retrieval set $\mathcal{D}$ of as little as 1,0001,0001 , 000 FPV images by letting agents navigate the environment and store images in $\mathcal{D}$. During training and evaluation, the retrieval-augmented agent can select dataset images from the scene it operates in. Additional details on dataset creation are provided in Appendix˜A].
@@
Implementation details – We base our agents on the DEBiT architecture by starting from the official codebase and weights provided by the authors [^2] https://github.com/naver/debit, and extend them as described in Section˜3. We consider two DEBiT backbones, base (DEBiT-b) with a binocular encoder with 55M parameters, and tiny (DEBiT-t), with a 17M-parameters encoder, to allow faster execution of ablation studies. The corresponding retrieval-augmented models are RANa-b and RANa-t. In Section˜5.2 we use the same agents and weights in zero-shot fashion for Instance-ImageNav and ObjectNav when the task goal is replaced with a retrieved image goal ([Uncaptioned image], [Uncaptioned image]). More details about the agent architecture are provided in Appendix˜B.
Model training – All models, retrieval-augmented or not, are trained with PPO up to 200M steps. Unless specified otherwise, the geometric FM $g$ and encoders $x$ and $l$ are loaded from DEBiT and kept frozen, as is DINOv2. For retrieval-augmented agents, the context encoder $c$ and policy network $\pi$ are learned from scratch. Compared to DEBiT, the GRU memory of RANa receives the additional context input $\tilde{\mathbf{c}}_{t}$, therefore we expand the DEBiT GRU input matrix to reach the required dimension and finetune it. Further details and ablations of this choice are reported in Appendix˜D. The reward definition is inspired by PointNav (Chattopadhyay et al., 2021) and ImageNav (Bono et al., 2024a) rewards and given as:
$r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$
where $K{=}10$, $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal, and slack cost $\lambda{=}0.01$ encourages efficiency.
Retrieval – We consider two context encoding approaches, Gumbel and Attention, and two retrieval mechanisms, static based on DINOv2, and dynamic based on DINOv2-Graph. The context size is $N{=}8$ in all experiments (see Appendix˜D for ablations). For dynamic variants requiring DINOv2 to be applied to each observation, we again leverage retrieval to speed up RL training significantly: during training, we replace observations $\mathbf{x}_{t}$ by their nearest dataset image, allowing to use pre-computed DINOv2 features.
Metrics – Navigation performance is evaluated by success rate (SR), i.e., the percentage of episodes terminated within a distance of <1{<}1< 1m from the goal after the agent has called the STOP action, and SPL (Anderson et al., 2018), i.e., SR weighted by the optimality of the path:
$\textit{SPL}=\frac{100}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$
where $S_{i}$ is a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
5.1 Retrieval improves navigation
Model Retrieval Context Enc. SR SPL
[Uncaptioned image] DEBiT-t (Bono et al., 2024a) - - 79.9 51.4
[Uncaptioned image] RANa-t static DINOv2 Gumbel 83.0 52.3
@@ 5.1 Retrieval improves navigation
[Uncaptioned image] RANa-t dynamic DINOv2-Graph Attention 84.7 63.7
[Uncaptioned image] DEBiT-b (Bono et al., 2024a) - - 83.4 56.8
[Uncaptioned image] RANa-b static DINOv2 Gumbel 88.3 57.6
@@
[Uncaptioned image] RANa-b dynamic DINOv2-Graph Attention 90.7 71.8
Table 2: Effect of retrieving context, RANa-t and -b models.
@@
Table˜2 shows results of the proposed RANa agents with retrieved context on the ImageNav task, exhibiting significant performance improvements across all settings compared to the DEBiT baselines [^3] To make baselines comparable, loaded checkpoints continued training for the same additional 200M steps, updating GRU and policy only.. For static context retrieval the Gumbel encoder shows better results, arguably because there is no inherent ordering in the context and the retrieved CroCo-based features contain strong information about the relative pose between context elements and observation. To simplify notation, from now on we label the combination of static context retrieval and Gumbel encoding static. The attention-based encoder works better for the dynamic graph-based context, where reasoning about the order of context elements becomes important. We label this graph-based context combined with attention encoding dynamic. It obtains the best results, with a huge boost in SPL (+15+15+ 15 and +12.3+12.3+ 12.3 points) compared to the DEBiT-b and DEBiT-t baselines, respectively.
Model Odom SR SPL Binocular encoder
OVRL (Yadav et al., 2023b) 54.2 27.0 Finetuned
VC1-ViT-L (Majumdar et al., 2023) 81.6 - Finetuned
OVRL-v2 (Yadav et al., 2023a) 82.0 58.7 Finetuned
ZSEL (Al-Halah et al., 2022) 29.2 21.6 Obs. & policy frozen, goal from scratch
ZSON (Majumdar et al., 2022) 36.9 28.0 Observation finetuned, goal frozen
FGPrompt (Sun et al., 2024a) 92.3 68.5 Trained from scratch
DEBiT-b no adapters (Bono et al., 2024a) 83.0 55.6 Frozen
RANa-b dynamic 90.7 71.8 Frozen
DEBiT-L w. adapters (Bono et al., 2024a) 94.0 71.7 Finetuned with adapters
RANa-L w. adapters dynamic 94.5 76.2 Finetuned with adapters
Table 3: ImageNav state-of-the-art results. this work sets the Habitat simulation parameter sliding=True, simplifying the task significantly and making it incomparable with other methods (cf. (Monaci et al., 2025)).
Comparison to ImageNav SOTA – In Table˜3 we compare RANa with state-of-the-art (SOTA) ImageNav models. These typically finetune (with a navigation loss) visual encoders based on CLIP (Majumdar et al., 2022), ViT (Yadav et al., 2023b; Majumdar et al., 2023; Yadav et al., 2023a) and geometric FMs  (Bono et al., 2024a) used for observation-goal comparison. Not only RANa considerably boosts the performance of the strong DEBiT-b model, but it also enhance the SOTA performance of DEBiT-Large (Bono et al., 2024a) finetuned with adapters (Chen et al., 2023). While SR is only marginally improved, arguably because it is already very high (949494) and remaining failure cases are not solvable using context, path efficiency (expressed by SPL) increases significantly, from 71.771.771.7 to 76.276.276.2 points.
[…]
DEBiT-t 27M 79.3 50.0 13.2 - - - 1.7 1.1 10.4
DEBiT-b 66M 83.0 55.6 25.2 - - - 1.7 1.1 22.4
RANa-t static 32M 83.0 52.3 24.4 0.1 0.1 0.4 1.7 1.2 20.9
RANa-t dynamic 34M 84.7 63.7 38.5 11.9 0.5 2.2 1.7 1.2 21.0
Table 4: Inference runtime of different model components, in ms, timed on one Nvidia H100-80G GPU. Retrieval overhead is negligible, most runtime is spent on image feature computations (DINOv2, CroCo), which increase with retrieval because of additional operations needed to select and compare retrieved elements.
Retrieval is efficientTable˜4 shows model size and runtime of different components of selected DEBiT and RANa models. Retrieval time is negligible and scales favorably with the dataset size thanks to efficient approximate nearest neighbor search libraries such as Faiss (Douze et al., 2024). Inference time is slightly higher for RANa because the encoder $c$ computes CroCo features between context images and observation and goal (while DEBiT only runs CroCo between observation and goal). Still, RANa-t is smaller and faster than DEBiT-b while achieving similar performance, suggesting that retrieval augmentation can be a frugal alternative to model scaling. The variant with dynamic context adds the compute time required to run DINOv2 on each observation, but its runtime remains in the same ballpark of other agents.
@@
100 dataset images suffice to improve performanceTable˜5 shows that the size of the retrieval database can be small, 100100100 images per scene suffice for the considered Gibson scenes, and performance plateaus at 111k images. Additional experiments for Instance-ImageNav and ObjectNav are presented in Appendix˜D.
DB size 100 1k 10k 20k
@@
Table 5: Impact of retrieval database size - SR (SPL) for RANa-t with static context using different database sizes, combinations of train and test conditions.
[…]
Model Retrieval Features SR SPL
[Uncaptioned image] DEBiT-t - 79.9 51.4
[Uncaptioned image] RANa-t static CroCo Geometric FM 83.0 52.3
[Uncaptioned image] RANa-t static hwRN18 79.1 48.2
Table 6: Impact of Geometric FM on context representation.
Context is importantTable˜7 studies the behavior of RANa-t models when ablating the context content: (i) in domain: expected retrieved elements, and (ii) random: random images from the scene dataset. Providing random context images results in a performance drop of both RANa static and dynamic to baseline levels. This result supports two key properties of our approach: the context is useful and RANa learns to exploit it, and at same time RANa can filter out irrelevant context elements and is robust to misleading random retrieval.
Model in domain random
RANa-t static 83.0 (52.3) 77.0 (47.1)
RANa-t dynamic 84.7 (63.7) 80.3 (50.2)
Table 7: Impact of retrieved context - SR (SPL) for RANa-t models on different test conditions.
Examples of retrieved contextFig.˜4 shows three types of context for $N=8$. Top 8 most similar dataset images to the goal in DINOv2 feature space (top) and top 8 MMR most similar images filtered by MMR (middle), are static context types, i.e.they are fixed along one episode and depend only on the goal image, shown on the right highlighted in red. DINOv2-Graph (bottom) is a dynamic context, recalculated at each step and depending on the observation (left, highlighted in blue).
The top 8 most similar images to the goal (top) are very similar to each other as well, and might not be very useful. Filtering images using MMR increases diversity and performance (cf. Appendix˜D for a qualitative analysis). The dynamic context (bottom) is the most informative, as it provides some form of “waypoints” from the current location to the goal – indeed this approach achieves the best navigation performance, as shown in Tables˜2 and 3.
Refer to captionStatic, top 8Static, top 8 MMRDynamic, DINOv2-Graph
Figure 4: Context types - top 8 most similar dataset images to the goal (red, on the right), top 8 MMR, most similar images filtered by MMR, and dynamic, the only one depending on the observation (blue, left) as well.